Find the measures of two complementary angles whose difference is 8
Answer by nerdybill(7384)

Find the measures of two complementary angles whose difference is 8
.
Sum of two complementary angles is 90 degrees.
.
Let x = one angle
then
90-x = second angle
.
our equation
(90-x) - x = 8
90-2x = 8
-2x = -82
x = 41 degrees
.
Second angle:
90-x = 90-41 = 49 degrees
.
answer: 41 and 49 degrees